Fermi National Laboratory
 

Accelerator Update for July 9, 2007 - July 11, 2007

 

Monday July 9

The day shift began with Operations monitoring the Tevatron (TeV) store 5543 (86.24E30) and the Recycle stash (89.5E10), with the Antiproton Source (Pbar) stacking (33.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI, MiniBooNE, and the Meson MTest experiment taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

An RF expert checked out MIRF11 and then brought it back on line at 8:24 AM.

The Neutron Therapy Facility (NTF) reported at 11:20 AM that they had finished treating their last patient for the day.

At 11:38 AM, a Linac Klystron RF station (KRF2) tripped off. It reset.

At 12:36 PM, two Booster RF stations (BRF13 & 14) tripped off when a vacuum pump (IPRF23) tripped off. Everything reset.

At 1:37 PM, the NuMI beam permit tripped off due to beam in the kicker gap. It reset.

Operations reported at 3:58 PM that the Recycler had stashed antiprotons three times during the day shift.

At 4:38 PM, a Booster power supply (GMPS) tripped off when the vacuum valves VB13 and VB17 closed. The valves would not open remotely. Operators reset them locally and turned on GMPS.

Operations disabled the MTest beam at 6 PM for the night.

At 6:47 PM, the 750keV line buncher tripped off. It reset.

At 8 PM, Operations reported that MIRF11 had tripped off many times for various reasons. Eventually, operators lowered the MI intensity.

At 8:07 PM, GMPS tripped off when the Booster vacuum valves closed for the first of three times over nine minutes. Operators reported that each trip was associated with a clap of thunder. After the last trip, operators noticed a vacuum gauge that was acting oddly (PG-16). They talked with an expert who suggested changing the gauge’s control board if the trips continued.

At 9:35 PM, LRF2 tripped off. It reset.

At 11:30 PM, the Pbar lithium lens and the sweeper magnet power supply tripped off. Operators contacted an expert.

At 11:43 PM, BRF3 tripped off. It reset.

Operations reported at 11:59 PM that the Recycler had stashed three times during the evening shift.

Tuesday July 10

The midnight shift began with Operations monitoring store 5543 (10.24E30) and the stash (296.3E10), with Pbar stacking (27.6E10), and with NuMI and MiniBooNE taking beam.

At midnight, a TeV expert began an end-of-store collimator study.

Operations turned the Pbar lithium lens and sweeper magnet back on and Pbar resumed stacking at 12:10 AM.

At 12:14 AM, MIRF11 tripped off for the first four times in about an hour. NuMI tripped off with each MIRF11 trip. Everything was reset, but after the last trip at 1:11 AM, operators bypassed the RF station.

The TeV study ended and Operations terminated store 5543 at 2 AM.

Operations put the TeV into shot setup at 2:13 AM.

At 4:16 AM, NuMI tripped off. It reset.

Operations established store 5544 at 4:31 AM with an initial luminosity of 186.2E30.

Operations reported at 4:54 AM that the H- Source’s current had dropped during the shot setup. Operators adjusted it back up to nominal values.

Operations established beam to MTest at 6:03 AM.

Operations reported at 8:10 AM that the Recycler had stashed three times during the midnight shift.

An RF expert brought MIRF11 back on line at 8:32 AM.

At 9:29 AM, Operations reported that MIRF11 had tripped off twice, taking NuMI down with it. Everything reset.

At 12:30 PM, a MTest quadrupole power supply (MT5Q2) tripped off on a magnet over temperature interlock indication. The magnet’s temperature had cooled off enough so that MTest could resume taking beam at 1:10 PM.

A MI expert began a collimator study at 3:58 PM.

Operations reported at 4 PM that the Recycler had stashed three times during the day shift.

Operations disabled the MTest beam at 6:01 PM.

At 7:12 PM, Operations reported that a Pbar Accumulator RF station (ARF4) had caused a very inefficient first stash during the last transfers of anitprotons. Operators informed a Pbar expert.

At 8:07 PM, a Booster vacuum gauge (PG-16) had again caused vacuum valves to close and GMPS to trip off. Operators replaced the gauge’s controller card.

At 9:16 PM, a Tevatron Electron Lens (TEL2) tripped off on a vacuum interlock. It reset.

Operations reported at 11:59 PM that the Recycler had stashed three times during the evening shift.

Wednesday July 11

The midnight shift began with Operations monitoring store 5544 and the stash (307.5E10), with Pbar stacking (11.2E10), and with NuMI and MiniBooNE taking beam.

At 2:39 AM, MIRF11 tripped off and took down NuMI. Everything reset.

At 2:54 AM, Operations reported problems with the last transfer of antiprotons from Pbar to the Recycler. Operators halted the stash and resumed stacking.

Operations turned off Booster and Linac at 5 AM in preparation of scheduled CDF shutdown for the day.

Operations had the Pbar rings ready for access by 5:23 AM.

Operations terminated store 5544 at 5:40 AM.

Operations had CDF and D0 ready for access by 5:49 AM.

Operations prepared NuMI and MiniBooNE for access at 6:29 AM.

Operations had Meson ready for access at by 6:38 AM.

Operations reported at 8:05 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to allow CDF to complete its detector work, which should take about 16 hours. The Run coordinator will allow other machine maintenance work as long as it doesn’t interfere with the Recycler stash.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/11/2007    email Fermilab